ok my teacher gave us words problmes to do and im stuck. Here is the problem: Bill is 3 times as old as Peter. Six years from now, Bill will be twice as old as Peter will be then. Find their present ages. Im stuck on what the formula is.
Found 2 solutions by stanbon, ankor@dixie-net.com:
Answer by stanbon(75887)

Bill is 3 times as old as Peter.
b = 3p
-------------------
Six years from now, Bill will be twice as old as Peter will be then.
(b+6) = 2(p+6)
---------------------------
Substitute for "b" and solve for "p":
(3p+6) = 2(p+6)
3p+6 = 2p+12
p = 6 (peter's age now)
---
Since b = 3p,
b = 18 (bill'e age now)
=============================
Cheers,
Stan H.


Find their present ages.

Answer by ankor@dixie-net.com(22740)

Bill is 3 times as old as Peter.
Six years from now, Bill will be twice as old as Peter will be then.
Find their present ages
:
Write an equation for each statement;
:
"Bill is 3 times as old as Peter."
b = 3p
:
" Six years from now, Bill will be twice as old as Peter will be then."
b + 6 = 2(p+6)
b + 6 = 2p + 12
b = 2p + 12 - 6
b = 2p + 6
:
From the 1st statement, replace b with 3p
3p = 2p + 6
3p - 2p = 6
p = 6 yrs is Peter's present age
:
I'll let you find Bill's age, check solutions in the 2nd statement